# Hermitian adjacency matrices with at most three distinct eigenvalues

### v1

@@ 4.1 $k=3$
Let $\gamma$ be a primitive third root of unity. Since the matrices $H_{\gamma}(D)$ and $H^{T}_{\gamma}(D)$ are similar we may assume, without loss of generality, that $\gamma=-\frac{1}{2}-i\frac{\sqrt{3}}{2}$. This implies that, for any oriented graph, $H_{\gamma}(D)=-H_{\omega}(D)$. So the $k=3$ variant of our problem is essentially identical to the $k=6$ case that we characterized in Section 2.
@@ 4.2 $k=4$
$\left(\begin{matrix}iI&0\\ 0&I\end{matrix}\right)\left(\begin{matrix}0&B\\ B^{T}&0\end{matrix}\right)\left(\begin{matrix}-iI&0\\ 0&I\end{matrix}\right)=i\left(\begin{matrix}0&B\\ -B^{T}&0\end{matrix}\right)=H.$
Where $H=H_{i}(D)$ is the Hermitian adjacency matrix of some oriented bipartite graph $D$. So $S$ and $H$ are similar, and $D$ has eigenvalues ±$\pm\theta$. The converse direction is very similar. ∎
Theorem 4.2 applies to many known signed graphs. Most notably, to an infinite family of 4-regular signed graphs (see [7]), and an infinite family of signed hypercubes, see [8].
It is possible that the assumption that $G$ is bipartite can be relaxed somewhat, but it cannot be dropped altogether: There is a signed $K_{6}$ with eigenvalues ±$\pm\sqrt{5}$, but we checked, with a computer, all orientations of $K_{6}$ and found that none of them yield a matrix $H_{i}(D)$ with two eigenvalues.
@@ 4.3 Other values of $k$
Now let $\sigma$ be a primitive $k$th root of unity. Certain aspects of our work in Section 2 remain applicable: Our first key result, Lemma 2.1, carries over nicely as long as $\mbox{Re}(\sigma)>0$.
@@
The calculations (2) and (3) from the proof of Lemma 2.1 remain valid. But now we have
$r\geq\frac{j^{\intercal}Hj}{j^{\intercal}j}=\frac{1}{n}\sum_{i>j}(H_{i,j}+H_{j,i})=\frac{2\mbox{Re}(\sigma)|E(D)|}{n}=2\mbox{Re}(\sigma)d,$
with equality if and only if $D$ is regular. Using (2) and (3) we can bound $s$ by
$s=\frac{mr}{m-n}=\frac{\left(\frac{2(n-m)d}{r}\right)}{m-n}=-2\left(\frac{d}{r}\right)\geq\frac{-1}{\mbox{Re}(\sigma)}.$
If we restrict to a primitive $k$th root of unity with maximal real part, then $\mbox{Re}(\sigma)$ is an increasing function of $k$ and the above bound becomes very strong.
@@
Let $k>8$ be an integer and let $\sigma=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$. The orientation of $K_{2}$ is the only connected oriented graph $D$ such that $H_{\sigma}(D)$ has exactly two distinct eigenvalues.
@@
First suppose $D$ is not a tournament. Then it contains an induced path on three vertices. Any orientation of the path on three vertices has $H_{\sigma}$-eigenvalues $\{-\sqrt{2},0,\sqrt{2}\}$, regardless of the value of $k$. So interlacing implies that the least eigenvalue of $H_{\sigma}(D)$ is at most $-\sqrt{2}$. But $\mbox{Re}(\sigma)=\cos(\frac{2\pi}{k})$ is an increasing function of $k$ whose value at $k=8$ is $\frac{1}{\sqrt{2}}$. So for any larger $k$, the bound from Corollary 4.3 implies $s>-\sqrt{2}$, a contradiction.
@@
$0=\operatorname{tr}(H_{\sigma}(D))=\sum_{i=1}^{n}\lambda_{i}\geq 2(n-1)\cos(\frac{2\pi}{k})-(n-2)\frac{1}{\cos(\frac{2\pi}{k})}.$
@@
So $r$ must have multiplicity 1 which implies $s$ has multiplicity $n-1$. The $H_{\sigma}$-adjacency matrix of an oriented $K_{2}$ has eigenvalues $\{1,-1\}$ and these eigenvalues must interlace the eigenvalues of $H_{\sigma}(D)$, but since $s$ has multiplicity $n-1$ this implies $s\geq-1\geq s$, so $s=-1$. But the $H_{\sigma}$-adjacency matrices of the two distinct orientations of the triangle each have an eigenvalue less than $-1$, so there are no induced triangles in our tournament and $n=2$. ∎
In the opposite extreme, one can use the fact that
$s\leq\frac{j^{\intercal}Hj}{j^{\intercal}j}=2\mbox{Re}(\sigma)d$
to get a strong upper bound on $r$ when $k$ is large and $\mbox{Re}(\sigma)$ is close to $-1$.
The intermediary choices, where $\mbox{Re}(\sigma)$ is close to zero, yield weaker bounds. This is most evident with $k=4$ where there exist oriented $n$-cubes which have $H_{i}$-eigenvalues ±$\pm\sqrt{n}$. When $\mbox{Re}(\sigma)$ is near zero we seem to have the best chance of finding large families of examples, but we do not yet know if there are infinite families when $k\neq 4$. So we conclude with a question.

### v2

@@ 4.1 $k=3$
Let $\gamma$ be a primitive third root of unity. Since the matrices $H_{\gamma}(D)$ and $H^{T}_{\gamma}(D)$ are similar we may assume, without loss of generality, that $\gamma=-\frac{1}{2}-i\frac{\sqrt{3}}{2}$. This implies that, for any oriented graph, $H_{\gamma}(D)=-H_{\omega}(D)$. So for oriented graphs the $k=3$ variant of our problem is essentially identical to the $k=6$ case that we characterized in Section 2.
@@
However, for mixed graphs the situation is very different. For each even $n$ there are non-trivial example of mixed orientations of $K_{3^{n}}$ with exactly two $H_{\gamma}$-eigenvalues. See Section 12.5 of [BCN89].
@@ 4.2 $k=4$
$\left(\begin{matrix}iI&0\\ 0&I\end{matrix}\right)\left(\begin{matrix}0&B\\ B^{T}&0\end{matrix}\right)\left(\begin{matrix}-iI&0\\ 0&I\end{matrix}\right)=i\left(\begin{matrix}0&B\\ -B^{T}&0\end{matrix}\right)=H,$
where $H=H_{i}(D)$ is the Hermitian adjacency matrix of some oriented bipartite graph $D$. So $S$ and $H$ are similar, and $D$ has two $H_{i}$-eigenvalues. The converse direction is very similar. ∎
Theorem 4.2 applies to many known signed graphs. Most notably, to an infinite family of 4-regular signed graphs (see [HTW19]), and an infinite family of signed hypercubes, see [HH19].
It is possible that the assumption that $G$ is bipartite can be relaxed somewhat, but it cannot be dropped altogether: There is a signed $K_{6}$ with eigenvalues $\pm\sqrt{5}$, but we checked, with a computer, all orientations of $K_{6}$ and found that none of them yield a matrix $H_{i}(D)$ with two eigenvalues.
@@ 4.3 Other values of $k$
Now let $\sigma$ be any primitive $k$th root of unity. Certain aspects of our work in Section 2 remain applicable: Our first key result, Lemma 2.1, carries over nicely as long as $\mbox{Re}(\sigma)>0$.
@@
The minimal polynomial of $H$ is still quadratic, so equations (2) and (3) from the proof of Lemma 2.1 remain valid. Now we have
$r\geq\frac{j^{\intercal}Hj}{j^{\intercal}j}=\frac{1}{n}\sum_{i>j}(H_{i,j}+H_{j,i})=\frac{2|E(D)|\mbox{Re}(\sigma)}{n}=\frac{(-nrs)\mbox{Re}(\sigma)}{n}$
with equality if and only if $D$ is regular. The bound on $s$ follows. ∎
If we restrict to the primitive $k$th root of unity $\sigma_{1}$ with maximal real part, then $\mbox{Re}({\sigma_{1}})$ is an increasing function of $k$ and the above bound becomes very strong.
@@
Let $k>8$ be an integer and let $\sigma_{1}=\cos(\frac{2\pi}{k})+i\sin(\frac{2\pi}{k})$. The orientation of $K_{2}$ is the only connected oriented graph $D$ such that $H_{\sigma_{1}}(D)$ has exactly two distinct eigenvalues.
@@
First suppose $D$ is not a tournament. Then it contains an induced path on three vertices. Any orientation of the path on three vertices has $H_{\sigma_{1}}$-eigenvalues $\{-\sqrt{2},0,\sqrt{2}\}$, regardless of the value of $k$. So interlacing implies that the least eigenvalue of $H_{\sigma_{1}}(D)$ is at most $\sqrt{2}$. But $\mbox{Re}({\sigma_{1}})=\cos(\frac{2\pi}{k})$ is an increasing function of $k$ whose value at $k=8$ is $\frac{1}{\sqrt{2}}$. So for any larger $k$, the bound from Corollary 4.3 implies $s>-\sqrt{2}$, a contradiction.
@@
$0=\operatorname{tr}(H_{\sigma_{1}}(D))=\sum_{i=1}^{n}\lambda_{i}\geq 2(n-1)\cos(\frac{2\pi}{k})-(n-2)\frac{1}{\cos(\frac{2\pi}{k})}.$
@@
So $r$ must have multiplicity 1 which implies $s$ has multiplicity n1n-1. The $H_{\sigma}$-adjacency matrix of an oriented $K_{2}$ has eigenvalues $\{1,-1\}$ and these eigenvalues must interlace the eigenvalues of $H_{\sigma_{1}}(D)$, but since $s$ has multiplicity n1n-1 this implies $s\geq-1\geq s$, so s=1s=-1. But the $H_{\sigma_{1}}$-adjacency matrices of the two distinct orientations of the triangle each have an eigenvalue less than 1-1, so there are no induced triangles in our tournament and $n=2$. ∎
In the opposite extreme, for a primitive $k$th root $\sigma_{2}$ with $\mbox{Re}(\sigma_{2})$ close to 1-1, one can use the fact that
$s\leq\frac{j^{\intercal}Hj}{j^{\intercal}j}=2\mbox{Re}(\sigma_{2})d$
to get a strong upper bound on $r$ when $k$ is large.
The intermediary choices, where $\mbox{Re}(\sigma)$ is close to zero, yield weaker bounds. This is most evident with $k=4$ where there exist oriented $n$-cubes which have $H_{i}$-eigenvalues $\pm\sqrt{n}$. When $\mbox{Re}(\sigma)$ is near zero we seem to have the best chance of finding large families of examples, but we do not yet know if there are infinite families when $k\neq 4$. So we conclude with a question.
@@
Conflict of Interest
Given his role as Associate Editor, Bojan Mohar had no involvement in the peer review of this article and had no access to information regarding its peer review. Full responsibility for the editorial process for this article was delegated to another journal editor.